OpenStudy (anonymous):

A Wheel rotated at 100 rpm. A circular body is placed on the wheel to reduce the angular velocity to 50rpm in 2s . The reduction in the angular velocity takes place in n revelutions of the wheel Then n=?

OpenStudy (anonymous):

Tell me if I'm right guys

OpenStudy (anonymous):

The answer should be 5/2

OpenStudy (anonymous):

@mahmit2012 HELP

OpenStudy (anonymous):

Use the equation \[\omega_f^2 - \omega_0^2 = 2\alpha \Delta \theta\] This equation works in radians and seconds but also in revolutions and minutes.... Here alpha is \[ \alpha = \frac{50-100}{2} rpm/s = -25 rpm/s = -1500 \frac{rpm}{m}\] So \[\Delta \theta = \frac{\omega_f^2 - \omega_0^2}{2\alpha} = \frac{50^2-100^2}{2\cdot -1500} = \frac{5}{2} \text{rotations}\] Just be sure to keep your units consistent. Dimensional analysis is very helpful in cases like these.

OpenStudy (anonymous):

thxxxx